OpenStudy (anonymous):

One number exceeds another by 26. The sum of the numbers is 64. What are the numbers?

OpenStudy (anonymous):

i got 19 and 45 ....

OpenStudy (anonymous):

64/2 = 32 +12 = 44

OpenStudy (anonymous):

THE NUMBERS ARE 'X' AND 'X+26' if you add the numbers the sum is 64 therefore x + x + 26 = 64 2x = 64 - 26 2x = 38 x = 19 x + 26 = 45 ;)